Consider the vector A⃗ with components Ax = -2.00, Ay= 6.00, the vector B⃗ with components Bx = 3.00, By= -1.00, and the vector  D⃗ =A⃗ −B⃗. Calculate the magnitude of the vector D⃗. (Express to three sig figs)
Answer #1

The vector A is, A= 4 i +4, j =(-2.00)i + (6.00); The vector B is, B=Bi +B, j =(3.00) { +(-1.00) | The new vector D is, D=A-B